OpenStudy (anonymous):

A restaurateur determines that the average daily revenue per table is given by R(x)=-1/4x+15 when there are x tables in the dining room. The average daily cost per table is $5.00.Determine an expression for the average daily profit per table. Enter your answer exactly in the form: P(X)=-a/bx+c

OpenStudy (campbell_st):

Profit = revenue - cost..... the C(x) = 5x then P(x) = R(x) - C(x) = -1/4x + 15 - 5x P(x) = -26/4 x + 15

OpenStudy (anonymous):

would it be -21/4x+15 since -1/4-5 is -21/4

OpenStudy (campbell_st):

yep